# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v1

@@ H.0.2 TopK SAE
The TopK SAE (Makhzani and Frey [2013], Gao et al. [2024]) directly controls the sparsity of the representation $\bm{z}$ by fixing it at $\|\bm{z}\|_{0}=K$, instead of imposing an explicit sparsity penalty $\mathcal{R}(\bm{z})$ in the loss function. The objective function for TopK SAE is:
@@ H.0.3 BatchTopK SAE
BatchTopK SAE (Bussmann et al. [2024]) is a modification of the TopK SAE. Instead of fixing sparsity like TopK, BatchTopK allows variable sparsity per input while fixing the mean sparsity over a batch at $K$. The objective function for BatchTopK SAE can equivalently be written as:
@@ H.0.4 JumpReLU SAE
JumpReLU SAE (Rajamanoharan et al. [2024]) is trained with the $L_{0}$ (pseudo-)norm regularizer. This leads to the following optimization problem:

### v2

@@ H.0.2 TopK SAE
The TopK SAE (makhzani2013k, gao2024scaling) directly controls the sparsity of the representation $\bm{z}$ by fixing it at $\|\bm{z}\|_{0}=K$, instead of imposing an explicit sparsity penalty $\mathcal{R}(\bm{z})$ in the loss function. The objective function for TopK SAE is:
@@ H.0.3 BatchTopK SAE
BatchTopK SAE (bussmann2024batchtopksparseautoencoders) is a modification of the TopK SAE. Instead of fixing sparsity like TopK, BatchTopK allows variable sparsity per input while fixing the mean sparsity over a batch at $K$. The objective function for BatchTopK SAE can equivalently be written as:
@@ H.0.4 JumpReLU SAE
JumpReLU SAE (rajamanoharan2024jumpingaheadimprovingreconstruction) is trained with the $L_{0}$ (pseudo-)norm regularizer. This leads to the following optimization problem: